how can u do this equation 0.7n-1.5+7.3n=14.5

Answer by checkley71(8403)

FIRST YOU NEED TO COMBINE THE COMMON TERMS:
.7N+1.5+7.3N=14.5 THUS
.7N+7.3N=14.5-1.5
8N=13
N=13/8 ANSWER.
PROOF
.7*13/8+1.5+7.3*13/8=14.5
9.1/8+1.5+94.9/8=14.5
9.1/8+94.9/8+1.5=14.5
104/8+1.5=14.5
13+1.5=14.5
14.5=14.5